# Counting Collatz Numbers

### v1

@@ 1 INTRODUCTION
$\pi(x)\geq cx^{H_{2}(1/\log_{2}3)},\ x\rightarrow+\infty,$
@@
2 A Congruence Equation
In this section we investigate an exponential congruence equation. We give a method to construct its solutions from free tuples.
@@ 2 A Congruence Equation
Lemma 2.2
Let $(u_{1},\cdots,u_{l})$ be an $l$-tuple of numbers. Then there is a unique solution $(v_{1},\cdots,v_{l})$ of the $(2,3)$-primary congruence equation of level $l$ such that
$\frac{v_{j}+1}{2}]=u_{j},\ j=1,\cdots,l.$
Proof. Firstly, if $l=1$, then it is easy to see that $v_{1}=2u_{1}$ is the unique solution of the $(2,3)$-primary level-11 congruence equation
$2^{x_{1}}\equiv 1({\rm mod}3)$
such that
$\frac{v_{1}+1}{2}]=u_{1}.$
such that
$\frac{v_{j}+1}{2}]=u_{j},\ j=1,\cdots,l-1.$
Then
$2^{v_{1}+\cdots+v_{l-1}}\equiv\sum_{j=0}^{l-1}3^{j}2^{v_{j+2}+\cdots+v_{l-1}}({\rm mod}3^{l-1}).$
Let
$a=3^{1-l}(2^{v_{1}+\cdots+v_{l-1}}-\sum_{j=0}^{l-1}3^{j}2^{v_{j+2}+\cdots+v_{l-1}}).$
Then $a\neq 0({\rm mod}3)$, and there is a unique binary digit $r_{l}$ such that
$2^{r_{l}}\equiv a({\rm mod}3).$
Let
$v_{l}=2u_{l}-r_{l}.$
Then
$2^{v_{l}}a\equiv 1({\rm mod}3).$
Let
$b=\frac{2^{v_{l}}a-1}{3}.$
Then
$2^{v_{1}+\cdots+v_{l}}-\sum_{j=0}^{l}3^{j}2^{v_{j+2}+\cdots+v_{l}}=3^{l}b\equiv 0({\rm mod}3^{l}),$
and
$\frac{v_{j}+1}{2}]=u_{j},\ j=1,\cdots,l.$
$\frac{v_{j}+1}{2}]=u_{j},\ j=1,\cdots,l.$
3 Collatz numbers
In this section relate Collatz numbers to solutions of the $(2,3)$-primary congruence equation.
@@ 3 Collatz numbers
Let $n>1$ be a Collatz number of level $l$, and
@@
then $n$ is a Collatz number of level $l$, and
@@
In particular, $(v_{1},v_{2},\cdots,v_{l-1})$ is a solution of the $(2,3)$-primary congruence equation of level $l$. By induction, $S(n)$ is a Collatz number of level (l1)(l-1), and
@@
and Therefore, $n$ is a Collatz number of level $l$, and
@@ 4 Bounds
$\omega(y,l)=\#\{(u_{1},\cdots,u_{l})\mid\sum_{j=1}^{l}u_{j}\leq y,(u_{1},\cdots,u_{l})\text{ is a tuple of numbers with }u_{1}>1\}.$
@@
$\sum_{j=1}^{l}u_{j}\leq\log_{4}x+l\log_{4}3.$
That is, $(u_{1},\cdots,u_{l})$ is an ordered $l$-partition with $u_{1}>1$ of a number $\leq\log_{4}x+l\log_{4}3.$ Let $(v_{1},\cdots,v_{l})$ be the unique solution of the $(2,3)$-primary congruence equation of level $l$ such that
$\frac{v_{j}+1}{2}]=u_{j},\ j=1,\cdots,l.$
@@
$B(v_{1},\cdots,v_{l})\leq 3^{-l}2^{v_{1}+\cdots+v_{l}}\leq 3^{-l}4^{u_{1}+\cdots+u_{l}}\leq x.$
@@
Lemma 4.4
Let $l=[\frac{\log_{2}x}{\log_{2}3}]$, and $n=[\log_{4}x+l\log_{4}3]-1$. There is a positive constant $c_{0}$ independent of $x$ such that, if $x$ is sufficiently large, then
$\log_{2}{n\choose l}\geq H_{2}(1/\log_{2}3)\log_{2}x-O(1),$
where $H_{2}(p)=-p\log_{2}p-(1-p)\log_{2}(1-p)$.
Proof. Notice that
$n-1=\log_{2}x+O(1),\text{ and }l/n=1/\log_{2}3+O(1/\log_{2}x).$
Thus, by Stirling’s formula,
$\log_{2}{n\choose l} \displaystyle\geq$ $nH_{2}(l/n)-O(1)$
$\geq$ $(\log_{2}x)H_{2}(1/\log_{2}3)-O(1).$
The lemma now follows.
Proof of Theorem 1.5. Let $l=[\frac{\log_{2}x}{\log_{2}3}]$, and $n=[\log_{4}x+l\log_{4}3]-1$. Then
@@
$\geq$ $2^{-c_{0}}x^{H_{2}(1/\log_{2}3)}.$
Choosing $c=2^{-c_{0}}$, we complete the proof of Theorem 1.5.

### v2

@@ 1 INTRODUCTION
$\pi(x)\geq(1+o(1))x^{\frac{1}{3}H_{2}(\frac{1}{2+\frac{2}{3}\log_{4}3})},\ x\rightarrow+\infty,$
@@
2 Entering Congruence Equations
In this section we relate Collatz numbers to a congruence equation.
@@ 2 Entering Congruence Equations
Let $n>1$ be a Collatz number prime to 33 of level $l$, and
@@
then $n$ is a Collatz number prime to 33 of level $l$, and
@@
In particular, $(v_{1},v_{2},\cdots,v_{l-1})$ is a solution of the $(2,3)$-primary congruence equation of level $l$. By induction, $S(n)$ is a Collatz number prime to 33 of level (l1)(l-1), and
@@
Therefore, $n$ is a Collatz number prime to 33 of level $l$, and
@@
3 Solving Congruence Equations
In this section we try to solve the congruence equation introduced in the last section. We give a method to construct its solutions from almost free tuples.
Lemma 3.1
Let $(u_{1},\cdots,u_{l})$ be an $l$-tuple of numbers prime to 33. Then there is a unique solution $(v_{1},\cdots,v_{l})$ of the $(2,3)$-primary congruence equation of level $l$ such that
$\frac{v_{j}+1}{6}]=u_{j}-1,\ j=1,\cdots,l.$
Proof. Firstly, if $l=1$, then it is easy to see that $v_{1}=6u_{1}-4$ is the unique solution of the $(2,3)$-primary level-11 congruence equation
[…]
such that
$\frac{v_{1}+1}{6}]=u_{1}-1.$
such that
$\frac{v_{j}+1}{6}]=u_{j}-1,\ j=1,\cdots,l-1.$
Then
$2^{v_{1}+\cdots+v_{l-1}}\equiv\sum_{j=0}^{l-2}3^{j}2^{v_{j+2}+\cdots+v_{l-1}}({\rm mod}3^{l-1}).$
Let
$a=3^{1-l}(2^{v_{1}+\cdots+v_{l-1}}-\sum_{j=0}^{l-2}3^{j}2^{v_{j+2}+\cdots+v_{l-1}}).$
Then $a\neq 0({\rm mod}3)$. Let $r_{l}$ be the unique binary digit such that
$2^{r_{l}}\equiv a({\rm mod}3),$
and let $w_{l}$ be the largest ternary digit such that
$4^{w_{l}}2^{r_{l}}a\not\equiv 1({\rm mod}3).$
Let
$v_{l}=6u_{1}-2w_{l}-r_{l}.$
Then
$2^{v_{l}}a\equiv 4^{u_{l}}\equiv 1({\rm mod}3),\text{ and }2^{v_{1}}a\not\equiv 1({\rm mod}3^{2}),$
and
$\frac{v_{j}+1}{6}]=u_{j}-1,\ j=1,\cdots,l.$
Let
$b=\frac{2^{v_{l}}a-1}{3}.$
Then
$2^{v_{1}+\cdots+v_{l}}-\sum_{j=0}^{l-1}3^{j}2^{v_{j+2}+\cdots+v_{l}}=3^{l}b\equiv 0({\rm mod}3^{l}),$
and
$2^{v_{1}+\cdots+v_{l}}-\sum_{j=0}^{l-1}3^{j}2^{v_{j+2}+\cdots+v_{l}}=3^{l}b\not\equiv 0({\rm mod}3^{l}),$
$\frac{v_{j}+1}{6}]=u_{j}-1,\ j=1,\cdots,l.$
@@ 4 Bounds
Lemma 4.1
Let $n,l\in\mathbb{N}$ with $\alpha n\leq l\leq n$, where $\alpha$ is a positive constant. Then
$\log_{2}{n\choose l}=H_{2}(l/n)n+O(1),\ n\rightarrow+\infty.$
Proof. This follows from the last lemma and Stirling’s formula.
@@
$\omega(y,l)=\#\{(u_{1},\cdots,u_{l})\in(\mathbb{N}+1)^{l}\mid\sum_{j=1}^{l}u_{j}\leq y\}.$
@@
$\sum_{j=1}^{l}u_{j}\leq\frac{1}{3}\log_{4}x+l(1+\frac{1}{3}\log_{4}3).$
That is, $(u_{1},\cdots,u_{l})$ is an ordered $l$-partition of a number $\leq\frac{1}{3}\log_{4}x+l(1+\frac{1}{3}\log_{4}3).$ Let $(v_{1},\cdots,v_{l})$ be the unique solution of the $(2,3)$-primary congruence equation of level $l$ such that
$\frac{v_{j}+1}{6}]=u_{j}-1,\ j=1,\cdots,l.$
@@
$B(v_{1},\cdots,v_{l})\leq 3^{-l}2^{v_{1}+\cdots+v_{l}}\leq(192)^{-l}4^{3(u_{1}+\cdots+u_{l})}\leq x.$
@@
Proof of Theorem 1.5. Let $l=[\frac{\log_{4}x}{3+1\log_{4}3}]$, and $n=[\frac{1}{3}\log_{4}x+l(1+\frac{1}{3}\log_{4}3)]$. Then
@@
$\geq$ $(1+o(1))x^{\frac{1}{3}H_{2}(\frac{1}{2+\frac{2}{3}\log_{4}3})}.$
This completes the proof of Theorem 1.5.